# Parametric Phase Modulation in Superconducting Circuits

### v1

@@ III.2 Parametric phase modulation
A transmon qubit under modulation exhibits various characteristics, requiring precise calibration for parametric phase modulation. When a tunable transmon is modulated, its time-averaged qubit frequency shifts due to the nonlinearity of transmon qubits. To track the frequency excursion, we perform three-tone spectroscopy experiments (see Appendix B for details). Notably, the amplitude-frequency response results in greater attenuation at higher parametric frequencies, which can be attributed to the hardware and fridge lines, commonly described by the transfer function (see Appendix C for details). Moreover, a tunable qubit driven by a parametric pulse can bridge the gap between two far-detuned qubits. Both $Q_{1}$ and $Q_{2}$ are tunable, and each qubit can independently be driven by a parametric pulse to induce effective two-qubit coupling (see Appendix D for details). These factors determine the calibrated operating points and pulse parameters (e.g., $\bar{\Phi}_{1},\tilde{\Phi}_{1},\omega_{p1},\bar{\Phi}_{2},\tilde{\Phi}_{2},\omega_{p2}$) for subsequent parametric phase modulation.
@@
Parametric phase modulation is a general method for adjusting coupling strength through phase. Both qubits are simultaneously subjected to two parametric pulses with the same sideband-resonant frequency, and the parametric relative phase modulates the effective two-qubit coupling strength. In the experiment, we demonstrate phase modulation for the first-order sideband couplings, tested at both sweet and off-sweet spots. We prepare the initial state $|10\rangle$ by applying a $\pi$-pulse on $Q_{1}$, followed by the simultaneous application of two parametric pulses to the qubits with the parameters below. At the sweet spots, we set the qubits’ bias to $\bar{\Phi}_{1}=\bar{\Phi}_{2}=0$. To control the first-order sideband coupling, we set $\tilde{\Phi}_{1}=0.08\Phi_{0}$, $\tilde{\Phi}_{2}=0.13\Phi_{0}$, and $\omega_{p1}/2\pi=\omega_{p2}/2\pi=70.8$ MHz. For the first-order sideband coupling at off-sweet spots, we use $\bar{\Phi}_{1}=0.064\Phi_{0}$, $\tilde{\Phi}_{1}=0.08\Phi_{0}$, $\bar{\Phi}_{2}=0.1025\Phi_{0}$, $\tilde{\Phi}_{2}=0.08\Phi_{0}$, and $\omega_{p1}/2\pi=\omega_{p2}/2\pi=178.64$ MHz. The relative parametric phase $\delta\phi_{p}$ is observed to control the period of population oscillations between the $|10\rangle$ and $|01\rangle$ states. Fig. 2 (a) presents the effective coupling strengths for the first-order ($n=1$) sideband interaction, which are extracted from these oscillations, along with corresponding theoretical fits based on Eq. (11). The range of coupling strengths achieved through this phase modulation is substantial, meeting typical requirements for implementing high-fidelity two-qubit gates or performing quantum simulations. To further demonstrate the versatility of our method, we have also applied and characterized this phase modulation technique for qubits interacting via parametric resonance (i.e., the zeroth-order sideband), with detailed results provided in Appendix F. A key advantage of our scheme is the suppression of frequency shifts when tuning the first-order coupling strength. This is explicitly demonstrated in Fig. 2(b), which numerically compares the frequency shifts induced by our phase modulation scheme against those from a conventional single-pulse amplitude modulation scheme. While the conventional method requires large adjustments to the parametric frequency to achieve a desired coupling strength, our scheme induces only negligible shifts across the entire tuning range. Furthermore, our scheme can achieve stronger coupling strengths, extending beyond the practical limits of the conventional amplitude-tuning approach [9]. Parametric phase modulation is a feasible method, requiring only the adjustment of the time-averaged qubit frequencies to achieve sideband resonance with the dual parametric pulses.
@@
In theory, the extremum of the phase-tunable coupling strength occurs at $\delta\phi_{p}=0$ for the first-order sideband coupling. However, two factors cause deviations from the extremum. First, the frequencies of the tunable transmon qubits depend non-linearly on the external flux, which leads to a nonlinear modulated frequency response and introduces additional phases beyond the linear region. Second, distortions in the actual pulses arise due to limitations of the microwave instruments and fridge lines, which inevitably introduce shifted local phases.
@@
Figure 2: Demonstration of phase-modulated coupling and its suppressed frequency shifts. (a) Phase-modulated coupling strength for the first-order ($n=1$) sideband, $2g_{\mathrm{phase}}^{1}/2\pi$, achieved with dual parametric pulses and demonstrated at sweet and off-sweet spots. Experimental data points are shown for the sweet spot (fuchsia circles) and the off-sweet spot (teal rhombuses). Corresponding dashed lines represent fits to these datasets using Eq. (11), rendered in distinct, high-contrast colors for clarity. The results from both operating conditions demonstrate that the relative parametric phase $\delta\phi_{p}$ effectively modulates the coupling strength. (b) Comparison of the induced qubit frequency shift required when tuning the coupling strength. Results from our phase modulation method ( fuchsia circles) are contrasted with those from conventional single-pulse amplitude modulation (teal rhombuses), highlighting the significant suppression of frequency shifts with our technique.
@@
We choose the parametric phase modulation using the first-order sideband coupling at the off-sweet spot for an illustration. The population oscillations between $|10\rangle$ and $|01\rangle$ are shown in Fig. 3(a), where the relative phase $\delta\phi_{p}$ modulates the coupling strength $2g_{\text{phase}}^{1}$. The coupling strength is largest at $\delta\phi_{p}=1.22\pi$ and smallest at $\delta\phi_{p}=0.18\pi$, as shown in Fig. 3(b). This demonstrates that the relative parametric phase $\delta\phi_{p}$ can modulate the coupling strength $2g_{\mathrm{phase}}^{1}$, while keeping the parametric amplitudes and frequencies fixed.
@@
Figure 3: Modulation of population oscillations between states $|10\rangle$ and $|01\rangle$ by the relative parametric phase $\delta\phi_{p}$. (a) Chevron pattern illustrating the population dynamics as a function of relative parametric phase $\delta\phi_{p}$ and evolution time. Oscillations at $\delta\phi_{p}=0.18\pi$ and $\delta\phi_{p}=1.22\pi$ (further detailed in (b)) are highlighted within the pattern. (b) Corresponding population oscillations versus evolution time at $\delta\phi_{p}=0.18\pi$ (fuchsia circles, fitted by brown solid line) and $\delta\phi_{p}=1.22\pi$ (teal stars, fitted by black dashed line). These traces demonstrate the phase-controlled modulation of the oscillation frequency, and thus the coupling strength.
@@
A spectroscopy experiment is a simple and effective technique for tracking the qubit frequency excursion. It also provides an intuitive way to determine the coupling strength from the gap in the avoided crossing. We conduct the spectroscopy experiment to observe the first-order sideband coupling at off-sweet spots, where the spectrum of $Q_{1}$ under parametric phase modulation is displayed in Fig. 4. This experiment is similar to the aforementioned three-tone spectroscopy for calibration, but with an additional parametric pulse applied to $Q_{2}$ at the same frequency called the four-tone spectroscopy, as illustrated in Fig. 4(a). An evident avoided crossing, corresponding to the effective coupling, is observed at the relative parametric phase $\delta\phi_{p}=0$, as seen in Fig. 4(b). The gap of the avoided crossing can be modulated by adjusting the parametric phase $\delta\phi_{p}$, while keeping other parameters fixed, as demonstrated in Fig. 4(c). This behavior is consistent with the results from population oscillations shown in Fig. 3(a). This spectroscopic measurement provides a frequency-domain confirmation of the coupling strength that is consistent with the time-domain population oscillations, offering a comprehensive validation of our phase-modulation technique.
@@
Figure 4: Spectroscopic observation of phase-modulated avoided crossings demonstrating tunable coupling. (a) Pulse sequence for the four-tone spectroscopy used to measure the $Q_{1}$ spectrum. (b) Measured spectrum of $Q_{1}$ under dual parametric pulses (on $Q_{1}$ and $Q_{2}$) at frequency $\omega_{p}/2\pi=70.8$ MHz, plotted as a function of varying parametric amplitude $\tilde{\Phi}_{1}$ on $Q_{1}$ (with $\tilde{\Phi}_{2}$ fixed at $0.13\Phi_{0}$). The inset shows a zoom-in of the avoided crossing corresponding to the first-order ($n=1$) sideband coupling, observed at $\tilde{\Phi}_{1}=0.08\Phi_{0}$. (c) The gap of the avoided crossing as a function of the relative parametric phase $\delta\phi_{p}$ between the dual parametric pulses. This demonstrates that $\delta\phi_{p}$ directly modulates the gap, i.e., the effective phase-tunable coupling strength $2g_{\text{phase}}^{1}/2\pi$.
@@
We explore the region where the flux-to-frequency transduction is nonlinear, i.e., qubits are near or parked at the DC sweet spot. Compared to the near-linear region, the non-linear region is harder to calibrate due to large frequency deviations under modulation. However, it holds practical significance due to its lower sensitivity to flux noise, as the dephasing rate depends on the power spectral density of the noise and the gradient of the qubit frequency with respect to flux variations $\partial\omega/\partial\Phi$ [49].
@@ IV Conclusion
We explore a wide range of tunable coupling strengths with the experimental parameters commonly used in quantum computation and simulation. We experimentally demonstrate the zeroth and first-order sideband coupling at both sweet and off-sweet spots. This approach offers a general method for modulating coupling strength, even for higher-order sidebands which owns high stability and scalability (see Appendix H for details).. Parametric phase modulation thus provides a new and versatile technique for adjusting coupling strength and extends the boundaries of parametric modulation. It is an additional, flexible tool for quantum simulations, and its application to reducing conditional phase or leakage errors in high-fidelity two-qubit gates is a promising avenue for future work.
@@ Acknowledgements.
This work was supported by the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0301702), NSFC (Grants No. U21A20436), NSF of Jiangsu Province (Grants No. BE2021015-1, and BK20232002), and Natural Science Foundation of Shandong Province (Grant No. ZR2023LZH002).
Appendix A Tunable coupling strength via a coupler
Figure 5: Qubit-qubit coupling strength $2g$ as a function of the coupler flux bias $\Phi_{c}$. (a) Measured population of the state $|10\rangle$, exhibiting oscillations as a function of evolution time and $\Phi_{c}$. (b) Extracted coupling strength $2g$, obtained by fitting the population oscillations from (a), plotted against $\Phi_{c}$. The vertical black dashed line indicates the selected working coupler flux, $\Phi_{c}=0.093\Phi_{0}$, with a coupling strength $2g/2\pi=21$ MHz used in the experiment.
@@ Appendix B Spectroscopy of parametrically modulated qubits
To demonstrate the effect of parametric flux pulses on qubits, we implement a three-tone spectroscopy experiment, using two microwave tones and a parametric flux tone, to capture the average response of qubit frequencies under parametric modulation [50, 42]. Unlike the two-tone spectroscopy experiment, an additional parametric flux pulse is applied to the qubit. We perform spectroscopy experiments on $Q_{1}$ at sweet $\bar{\Phi}_{1}=0$ and off-sweet $\bar{\Phi}_{1}=0.064\Phi_{0}$ spots, with fixed parametric frequencies of $\omega_{p1}/2\pi=79.2$ MHz and $\omega_{p1}/2\pi=181.2$ MHz, respectively, as shown in Fig. 6. We observe that the time-averaged qubit frequencies decrease as the parametric amplitudes increase with the fixed parametric frequency, which arises from the nonlinearity of the transmon qubit to the external flux. When the parametric amplitude approaches zero, the spectrum shows only a peak, representing the DC component of the qubit’s modulated frequency. However, as the parametric amplitude increases, harmonic frequency peaks appear, corresponding to the sideband frequencies of the modulated qubit. At off-sweet spots, these sidebands follow the pattern $\bar{\omega}_{1}+n\omega_{p1}$, and at sweet spots, they follow $\bar{\omega}_{1}+2n\omega_{p1}$, where $n\in\mathbb{Z}$.
This relationship between the time-averaged frequencies and parametric amplitudes provides insights into the parametric sideband-resonant conditions of two qubits and facilitates further calibration. At sweet spots, the first-order sideband frequencies $\bar{\omega}_{1}\pm\omega_{p1}$ theoretically disappear, which contrasts with the behavior at off-sweet spots. This is because modulated qubits undergo double cycles compared to the parametric flux pulses at sweet spots [10]. As shown in Fig. 6 (a), the first-order sidebands are nearly invisible because of the small flux deviations at sweet spots. The parametric amplitude $\tilde{\Phi}_{1}$ (related to the frequency excursion $\epsilon_{p1}$) determines the effective coupling strength $g_{\mathrm{eff}}^{n}$ of different sidebands via Bessel functions, as describe in Eq. (8). Consequently, variations in $\tilde{\Phi}_{1}$ directly impact the observed prominence and intensity of these sidebands in the spectrum shown in Fig. 6. An interesting phenomenon is that the linewidth of the modulated qubit broadens with larger amplitudes due to power broadening [50].

### v2

@@ III.1 Chip parameters
$\Phi_{1}(t)=\bar{\Phi}_{1}+\tilde{\Phi}_{1}\cos(\omega_{p1}t+\phi_{p1})$
and
$\Phi_{2}(t)=\bar{\Phi}_{2}+\tilde{\Phi}_{2}\cos(\omega_{p2}t+\phi_{p2}),$
are applied to the two qubits. Additionally, a static dc flux bias $\Phi_{c}=0.093\Phi_{0}$ is used to bias the coupler to achieve an appropriate two-qubit coupling strength $2g/2\pi=21$ MHz (see Appendix D for details), which depends on the frequencies and fixed coupling strength of this system [19].
@@ III.2 Parametric phase modulation
A transmon qubit under modulation exhibits various characteristics, requiring precise calibration for parametric phase modulation. When a tunable transmon is modulated, its time-averaged qubit frequency shifts due to the nonlinearity of transmon qubits. To track the frequency excursion, we perform three-tone spectroscopy experiments (see Appendix B for details). Notably, the amplitude-frequency response results in greater attenuation at higher parametric frequencies, which can be attributed to the hardware and fridge lines, commonly described by the transfer function (see Appendix E for details). Moreover, a tunable qubit driven by a parametric pulse can bridge the gap between two far-detuned qubits. Both $Q_{1}$ and $Q_{2}$ are tunable, and each qubit can be driven independently by a parametric pulse to induce effective two-qubit coupling (see Appendix F for details). These factors determine the calibrated operating points and pulse parameters (e.g., $\bar{\Phi}_{1},\tilde{\Phi}_{1},\omega_{p1},\bar{\Phi}_{2},\tilde{\Phi}_{2},\omega_{p2}$) for subsequent parametric phase modulation.
Parametric phase modulation is a general method for adjusting coupling strength through phase. Both qubits are simultaneously subjected to two parametric pulses with the same sideband-resonant frequency, and the parametric relative phase modulates the effective two-qubit coupling strength. In the experiment, we demonstrate phase modulation for the first-order sideband couplings, tested at both sweet and off-sweet spots. We prepare the initial state $|10\rangle$ by applying a $\pi$ pulse on $Q_{1}$, followed by the simultaneous application of two parametric pulses to the qubits with the parameters below. At the sweet spots, we set the qubits’ bias to $\bar{\Phi}_{1}=\bar{\Phi}_{2}=0$. To control the first-order sideband coupling, we set $\tilde{\Phi}_{1}=0.08\Phi_{0}$, $\tilde{\Phi}_{2}=0.13\Phi_{0}$, and $\omega_{p1}/2\pi=\omega_{p2}/2\pi=70.8$ MHz. For the first-order sideband coupling at off-sweet spots, we use $\bar{\Phi}_{1}=0.064\Phi_{0}$, $\tilde{\Phi}_{1}=0.08\Phi_{0}$, $\bar{\Phi}_{2}=0.1025\Phi_{0}$, $\tilde{\Phi}_{2}=0.08\Phi_{0}$, and $\omega_{p1}/2\pi=\omega_{p2}/2\pi=178.64$ MHz. The relative parametric phase $\delta\phi_{p}$ is observed to control the period of population oscillations between the $|10\rangle$ and $|01\rangle$ states.
@@
Fig. 2 (a) presents the effective coupling strengths for the first-order ($n=1$) sideband interaction, which are extracted from these oscillations, along with corresponding theoretical fits based on Eq. (11). The range of coupling strengths achieved through this phase modulation is substantial, meeting typical requirements for implementing high-fidelity two-qubit gates or performing quantum simulations. To further demonstrate the versatility of our method, we have also applied and characterized this phase-modulation technique for qubits interacting via parametric resonance (i.e., the zeroth-order sideband), with detailed results provided in Appendix G. A key advantage of our scheme is the suppression of frequency shifts when tuning the first-order coupling strength. This is explicitly demonstrated in Fig. 2(b), which numerically compares the frequency shifts induced by our phase-modulation scheme against those from a conventional single-pulse amplitude-modulation scheme. While the conventional method requires large adjustments to the parametric frequency to achieve a desired coupling strength, our scheme induces only negligible shifts across the entire tuning range. Furthermore, our scheme can achieve stronger coupling strengths, extending beyond the practical limits of the conventional amplitude-tuning approach [9]. Parametric phase modulation is a feasible method, requiring only the adjustment of the time-averaged qubit frequencies to achieve sideband resonance with dual parametric pulses.
@@
In theory, the extremum of the phase-tunable coupling strength occurs at $\delta\phi_{p}=0$ for the first-order sideband coupling. However, two factors cause deviations from the extremum. First, the frequencies of the tunable transmon qubits depend nonlinearly on the external flux, which leads to a nonlinear modulated frequency response and introduces additional phases beyond the linear region. Second, distortions in the actual pulses arise due to limitations of the microwave instruments and fridge lines, which inevitably introduce shifted local phases.
@@
Figure 2: Demonstration of phase-modulated coupling and its suppressed frequency shifts. (a) Phase-modulated coupling strength for the first-order ($n=1$) sideband, $2g_{\mathrm{phase}}^{1}/2\pi$, achieved with dual parametric pulses and demonstrated at sweet and off-sweet spots. Experimental data points are shown for the sweet spot (fuchsia circles) and the off-sweet spot (teal rhombuses). Corresponding dashed lines represent fits to these datasets using Eq. (11), rendered in distinct, high-contrast colors for clarity. The results from both operating conditions demonstrate that the relative parametric phase $\delta\phi_{p}$ effectively modulates the coupling strength. (b) Comparison of the induced qubit frequency shift required when tuning the coupling strength. Results from our phase-modulation method (fuchsia circles) are contrasted with those from conventional single-pulse amplitude modulation (teal rhombuses), highlighting the significant suppression of frequency shifts with our technique.
@@
We choose parametric phase modulation using the first-order sideband coupling at the off-sweet spot for an illustration. The population oscillations between $|10\rangle$ and $|01\rangle$ are shown in Fig. 3(a), where the relative phase $\delta\phi_{p}$ modulates the coupling strength $2g_{\text{phase}}^{1}$. The coupling strength is largest at $\delta\phi_{p}=1.22\pi$ and smallest at $\delta\phi_{p}=0.18\pi$, as shown in Fig. 3(b). This demonstrates that the relative parametric phase $\delta\phi_{p}$ can modulate the coupling strength $2g_{\mathrm{phase}}^{1}$, while keeping the parametric amplitudes and frequencies fixed.
@@
Figure 3: Modulation of population oscillations between states $|10\rangle$ and $|01\rangle$ by the relative parametric phase $\delta\phi_{p}$. (a) Chevron pattern illustrating the population dynamics as a function of relative parametric phase $\delta\phi_{p}$ and evolution time. Oscillations at $\delta\phi_{p}=0.18\pi$ and $\delta\phi_{p}=1.22\pi$ (further detailed in panel (b)) are highlighted within the pattern. (b) Corresponding population oscillations versus evolution time at $\delta\phi_{p}=0.18\pi$ (fuchsia circles, fitted by brown solid line) and $\delta\phi_{p}=1.22\pi$ (teal stars, fitted by black dashed line). These traces demonstrate the phase-controlled modulation of the oscillation frequency, and thus the coupling strength.
@@
A spectroscopy experiment is a simple and effective technique for tracking the qubit frequency excursion. It also provides an intuitive way to determine the coupling strength from the gap in the avoided crossing. We conduct the spectroscopy experiment to observe the first-order sideband coupling at off-sweet spots, where the spectrum of $Q_{1}$ under parametric phase modulation is displayed in Fig. 4. This experiment is similar to the aforementioned three-tone spectroscopy for calibration, but with an additional parametric pulse applied to $Q_{2}$ at the same frequency called four-tone spectroscopy, as illustrated in Fig. 4(a). An evident avoided crossing, corresponding to the effective coupling, is observed at the relative parametric phase $\delta\phi_{p}=0$, as seen in Fig. 4(b). The gap of the avoided crossing can be modulated by adjusting the parametric phase $\delta\phi_{p}$, while keeping other parameters fixed, as demonstrated in Fig. 4(c). This behavior is consistent with the results from population oscillations shown in Fig. 3(a). This spectroscopic measurement provides a frequency-domain confirmation of the coupling strength that is consistent with the time-domain population oscillations, offering a comprehensive validation of our phase-modulation technique.
@@
Figure 4: Spectroscopic observation of phase-modulated avoided crossings demonstrating tunable coupling. (a) Pulse sequence for the four-tone spectroscopy used to measure the $Q_{1}$ spectrum. (b) Measured spectrum of $Q_{1}$ under dual parametric pulses (on $Q_{1}$ and $Q_{2}$) at frequency $\omega_{p}/2\pi=70.8$ MHz, plotted as a function of varying parametric amplitude $\tilde{\Phi}_{1}$ on $Q_{1}$ (with $\tilde{\Phi}_{2}$ fixed at $0.13\Phi_{0}$). The inset shows an enlarged view of the avoided crossing corresponding to the first-order ($n=1$) sideband coupling, observed at $\tilde{\Phi}_{1}=0.08\Phi_{0}$. (c) The gap of the avoided crossing as a function of the relative parametric phase $\delta\phi_{p}$ between the dual parametric pulses. This demonstrates that $\delta\phi_{p}$ directly modulates the gap, i.e., the effective phase-tunable coupling strength $2g_{\text{phase}}^{1}/2\pi$.
@@
We explore the region where the flux-to-frequency transduction is nonlinear, i.e., qubits are near to or parked at the dc sweet spot. Compared to the near-linear region, the nonlinear region is harder to calibrate due to large frequency deviations under modulation. However, it holds practical significance due to its lower sensitivity to flux noise, as the dephasing rate depends on the power spectral density of the noise and the gradient of the qubit frequency with respect to flux variations $\partial\omega/\partial\Phi$ [49].
@@ IV Conclusion
We explore a wide range of tunable coupling strengths with the experimental parameters commonly used in quantum computation and simulation. We experimentally demonstrate the zeroth- and first-order sideband coupling at both sweet and off-sweet spots. This approach offers a general method for modulating coupling strength, even for higher-order sidebands, which possesses high stability and scalability (see Appendix H for details). Parametric phase modulation thus provides a new and versatile technique for adjusting coupling strength and extends the boundaries of parametric modulation. It is an additional, flexible tool for quantum simulations, and its application to reducing conditional phase or leakage errors in high-fidelity two-qubit gates is a promising avenue for future work.
DATA AVAILABILITY
The data that support the findings of this paper are openly available at [50].
@@ Acknowledgements.
This work was supported by the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0301702), the NSFC (Grant No. U21A20436), the NSF of Jiangsu Province (Grants No. BE2021015-1 and No. BK20232002), the Natural Science Foundation of Jiangsu Province (Grant No. BK20233001), and the Natural Science Foundation of Shandong Province (Grant No. ZR2023LZH002).
Appendix A Coherence under parametric phase modulation
Coherence is central to high-fidelity operations. For qubits under parametric modulation, coherence is influenced by both their intrinsic properties and the surrounding environment. The flux-noise spectra of qubits determine their dephasing rates, while modulation can sometimes increase dephasing due to effects like multiplicative $1/f$ noise. Beyond flux noise, the qubit’s environment, particularly two-level system (TLS) defects, is a common source of decoherence. Theoretical work has shown that frequency modulation can help stabilize relaxation rates and mitigate specific dephasing mechanisms arising from TLSs [31, 32]. The stability and quality of qubits can also be affected by slow periodic frequency modulation [33].
For the flux noise, there are powerful mitigation strategies. For instance, operating at specific modulation parameters can create “dynamical sweet spots” that are insensitive to $1/f$ flux noise [24, 14], a technique that has been successfully used to demonstrate high-fidelity controlled-Z (CZ) gates [26]. The dephasing rate $\Gamma_{\phi,1/f}$ under modulation can be described as
$\Gamma_{\phi,1/f}=\lambda\left|\frac{\partial\bar{\omega}}{\partial\tilde{\Phi}}\right|A_{\rm{ac},1/f},$ (12)
where $\lambda$ and $A_{\rm{ac},1/f}$ represent the noise parameter and amplitude due to $1/f$ noise [24]. This concept can be extended. For example, two-tone parametric modulation can create a continuum of dynamical sweet spots, expanding the range of flux-noise-robust operating frequencies [25, 29]. The benefits of parametric drives have also been explored for enhancing coherence in bichromatically driven Floquet qubits [30]. Parametric modulation can also be applied to fluxonium qubits to protect qubits from $1/f$ noise and enhance coherence times [28, 27]. Indeed, parametric modulation can be viewed as a continuous version of the dynamical decoupling scheme to realize high-fidelity CZ gates which can greatly reduce qubit dephasing [23].
Coherence degradation is a known challenge for some hardware-based tunable coupler schemes [51, 52]. Such degradation often occurs when the system is tuned into a highly hybridized regime, where the coherence of the qubits can be limited by that of the coupler which can be explained using a two-spin Hamiltonian [53] or by the resulting dressed states becoming more sensitive to flux noise. The effective decoherence rates in such systems can be modeled by the participation ratios of the uncoupled modes [54].
In our experiment, we have designed our scheme to specifically enhance coherence. We choose two qubits with a coupler in our chip, and other irrelevant couplers are biased to idle points with zero couplings. The chosen hardware coupler is biased to a fixed operating point that maintains a sufficient effective coupling strength while ensuring the qubits are only weakly hybridized with it. Crucially, our working points for the qubits are at or near to their dc sweet spots, leveraging their inherent insensitivity to flux noise [49]. Consequently, the dual parametric drives, which are simultaneously applied to the two adjacent qubits, are designed to inherit the robustness of parametric modulation against pulse distortions and noise. The introduction of a second, synchronous drive primarily adds a new degree of control—the relative phase—without fundamentally altering the underlying noise-protection mechanisms.
@@ Appendix B Spectroscopy of parametrically modulated qubits
To demonstrate the effect of parametric flux pulses on qubits, we implement a three-tone spectroscopy experiment, using two microwave tones and a parametric flux tone, to capture the average response of qubit frequencies under parametric modulation [55, 42]. Unlike the two-tone spectroscopy experiment, an additional parametric flux pulse is applied to the qubit. We perform spectroscopy experiments on $Q_{1}$ at sweet spot, $\bar{\Phi}_{1}=0$, and off-sweet spot, $\bar{\Phi}_{1}=0.064\Phi_{0}$, with fixed parametric frequencies of $\omega_{p1}/2\pi=79.2$ MHz and $181.2$ MHz, respectively, as shown in Fig. 5. We observe that the time-averaged qubit frequencies decrease as the parametric amplitudes increase with the fixed parametric frequency, which arises from the nonlinearity of the transmon qubit to the external flux. When the parametric amplitude approaches zero, the spectrum shows only a peak, representing the dc component of the qubit’s modulated frequency. However, as the parametric amplitude increases, harmonic frequency peaks appear, corresponding to the sideband frequencies of the modulated qubit. At off-sweet spots, these sidebands follow the pattern $\bar{\omega}_{1}+n\omega_{p1}$, and at sweet spots, they follow $\bar{\omega}_{1}+2n\omega_{p1}$, where $n\in\mathbb{Z}$.
This relationship between the time-averaged frequencies and parametric amplitudes provides insights into the parametric sideband-resonant conditions of two qubits and facilitates further calibration. At sweet spots, the first-order sideband frequencies $\bar{\omega}_{1}\pm\omega_{p1}$ theoretically disappear, which contrasts with the behavior at off-sweet spots. This is because modulated qubits undergo double cycles compared to the parametric flux pulses at sweet spots [10]. As shown in Fig. 5 (a), the first-order sidebands are nearly invisible because of the small flux deviations at sweet spots. The parametric amplitude $\tilde{\Phi}_{1}$ (related to the frequency excursion $\epsilon_{p1}$) determines the effective coupling strength $g_{\mathrm{eff}}^{n}$ of different sidebands via Bessel functions, as described in Eq. (8). Consequently, variations in $\tilde{\Phi}_{1}$ directly impact the observed prominence and intensity of these sidebands in the spectrum shown in Fig. 5. An interesting phenomenon is that the linewidth of the modulated qubit broadens with larger amplitudes due to power broadening [55].
@@
Appendix C Numerical comparison between the Taylor expansion and the Fourier series
Using the power-reduction formulas, we can rewrite the Taylor expansion (13) as
$\begin{split}\omega(t)&=\left[\omega(\bar{\Phi})+\sum_{n=1}^{\infty}\frac{\tilde{\Phi}^{n}}{2^{n}n!}\left.\frac{\partial^{n}\omega}{\partial\Phi^{n}}\right|_{\bar{\Phi}}\delta_{{n\bmod 2},0}\binom{n}{\lfloor{n}/{2}\rfloor}\right]\\ &+\sum_{n=1}^{\infty}\frac{2\tilde{\Phi}^{n}}{2^{n}n!}\left.\frac{\partial^{n}\omega}{\partial\Phi^{n}}\right|_{\bar{\Phi}}\sum_{k=0}^{\lfloor{(n-1)}/{2}\rfloor}\binom{n}{k}\cos((n-2k)\omega_{p}t).\end{split}$ (14)
As the expansion order $n$ increases, more harmonics appear, improving accuracy. We choose typical parameters $\bar{\Phi}=0,~\tilde{\Phi}=0.4\Phi_{0}$ at sweet spots and $\bar{\Phi}=0.15\Phi_{0},~\tilde{\Phi}=0.3\Phi_{0}$ at off-sweet spots as examples.
Figure 6: Comparison of Taylor expansion and Fourier series for calculating components of the parametrically modulated qubit frequency. (a) The dc frequency shift of the qubit under parametric modulation, shown as a function of the expansion order $n$. Results are presented for modulation at the sweet spot ($\bar{\Phi}=0,\tilde{\Phi}=0.4\Phi_{0}$; teal lines) and an off-sweet spot ($\bar{\Phi}=0.15\Phi_{0},\tilde{\Phi}=0.3\Phi_{0}$; fuchsia lines). Solid lines represent calculations from Taylor expansion, while dashed lines show the corresponding Fourier-series components. (b) Amplitude of the dominant ac harmonic component of the modulated qubit frequency versus expansion order $n$. For the sweet spot (teal lines), this is the second harmonic (frequency $2\omega_{p}$ component); for the off-sweet spot (fuchsia lines), it is the first harmonic (frequency $\omega_{p}$ component). Taylor-expansion results (solid lines) are shown approaching the Fourier-series values (dashed lines) as $n$ increases. All calculations use transmon qubit parameters $E_{C}/2\pi=240$ MHz and $E_{J1}/2\pi=E_{J2}/2\pi=8.286$ GHz.
As shown in Fig. 6 (a), the dc shifts depend on the even-order terms of the Taylor expansion. The main shift can be approximated by the second-order ($n=2$) term. As the Taylor order $n$ increases, the dc component approaches the zeroth order of the Fourier series, corresponding to the time-averaged frequency for one period. The dc component results from the nonlinearity in transmon qubits and can also be derived using the charge and flux operators, as shown in Ref. [42], rather than the analytical equation provided above.
As shown in Fig. 6 (b), we respectively estimate the second harmonic at the sweet spot and the first harmonic at the off-sweet spot, which are used to theoretically determine the parametric coupling strength at weak parametric modulation amplitudes. The main harmonic components are obtained by considering the Taylor expansion at order $n=2$ (the sweet spot) and at order $n=3$ (the off-sweet spot). At sweet spots, the qubit frequency oscillates only at even harmonics of the parametric frequency since all odd-order derivatives vanish. No significant difference is observed for Taylor-expansion orders $n>6$ at sweet and off-sweet spots.
Figure 7: Average deviations of Taylor-expansion and Fourier-series approximations from an analytical solution for the parametrically modulated qubit frequency $\omega(t)$. (a) Average deviation over one modulation period from the analytical solution for calculations at the sweet spot, plotted as a function of increasing expansion order $n$. (b) Corresponding average deviation for calculations at the off-sweet spot. In both panels (a) and (b), results for the Fourier series are shown as solid lines with circle markers, while results for the Taylor expansion are shown as solid lines with triangle-up markers. Insets: Time evolution of the modulated qubit frequency, $\omega(t)$, over one period, comparing the analytical solution (black solid line) with approximations from Fourier series (fuchsia dashed line) and Taylor expansion (teal dotted line) at expansion orders $n=3$ and $n=10$. All physical and modulation parameters are identical to those used in Fig. 6.
Figure 8: Average deviations of Taylor-expansion and Fourier-series approximations (at a fixed expansion order of $n=3$) from an analytical solution for the parametrically modulated qubit frequency, shown as the parametric amplitude $\tilde{\Phi}$ increases. Solid and dashed lines represent calculations for the qubit at its sweet and off-sweet spots, respectively. Fuchsia lines correspond to the Fourier series approximation, and teal lines correspond to the Taylor expansion approximation. All other physical and modulation parameters are identical to those specified in Fig. 6.
The time-dependent frequency of qubits under parametric modulation can be approximated using both the Taylor expansion and the Fourier series. The average deviations over a period between these methods and the analytical equation decrease as the expansion order $n$ increases at sweet and off-sweet spots, as shown in Fig. 7. The insets show the time evolution of qubit frequencies over one period at orders $n=3$ and $n=10$. At $n=3$, the Taylor expansion slightly deviates, while no visible deviations occur at higher orders. At sweet spots, the Fourier spectrum exhibits a stepped shape due to the vanishing of odd Fourier coefficients [9]. Average deviations are lower at sweet spots compared to off-sweet spots, due to the symmetry of flux-to-frequency transduction. Both methods provide good approximations of qubit time-dependent frequencies and coupling-strength calculations for weak parametric amplitudes, but the accuracy decreases as the parametric amplitude increases, as shown in Fig. 8. At sweet spots, both methods yield lower deviations due to the symmetry of flux-to-frequency transduction, which eliminates odd harmonics.
Appendix D Tunable coupling strength via a coupler
Figure 9: Qubit-qubit coupling strength $2g$ as a function of the coupler flux bias $\Phi_{c}$. (a) Measured population of the state $|10\rangle$, exhibiting oscillations as a function of evolution time and $\Phi_{c}$. (b) Extracted coupling strength $2g$, obtained by fitting the population oscillations from panel (a), plotted against $\Phi_{c}$. The vertical black dashed line indicates the selected working coupler flux, $\Phi_{c}=0.093\Phi_{0}$, with a coupling strength $2g/2\pi=21$ MHz used in the experiment.
Appendix E Flux pulse transfer function